alexandria:

Ivanova has a healthcare plan with $25 co-pays to see doctors, $10 co-pays for prescription drugs, and an annual deductible of $1,500. In February she has a broken ankle which costs $2,300 in hospital treatment. As a follow-up in March she sees three doctors who bill $657, $134, and $232 and buys two prescriptions which cost $12 and $27. What is Ivanova's share of the costs for March?

Extrinix:

Ivanova's plan covers the following: - Doctors: $25 copay - Prescriptions: $10 copay - Deductible: $1500 February - $2300 hospital treatment March - Doctor 1: billed $657, she pays $25, after deductible [+$25] - Doctor 2: billed $134, she pays $25, after deductible [+$25] - Doctor 3: billed $232, she pays $25, after deductible [+$25] - Prescription: cost $12, $10 copay [+$10] - Prescription: cost $27, $10 copay [+$10] As we don't consider the February, given we are looking at the share of cost for March, we can see that she visits 3 doctors (\(3\times 25\)), and gets 2 prescriptions (\(2\times 10\)). Therefore we can conclude the amount she pays is as follows: \(3(25)+2(10)\) I will let you solve that simple math problem to find out how much she pays in all. :)

Extrinix:

I should say ^ pays for that month as that is more accurate.

Aliciaa:

Yh wht @extrinix said. It's correct but there is no point in me sending the same thing when he alr said it lol!! Good Job!

marienne:

tht right

whypyt:

1842 I think ,hope this helps :D

forgetmylife:

Josh is right

GRIZZLEY19:

Ivanova's share of the costs for March is $795, which includes $75 in doctor co-pays, $20 in prescription co-pays, and $700 toward her deductible.

toga:

@aliciaa wrote:
Yh wht @extrinix said. It's correct but there is no point in me sending the same thing when he alr said it lol!! Good Job!
yea its @magaback want do you think?

MAGABACK:

@extrinix wrote:
Ivanova's plan covers the following: - Doctors: $25 copay - Prescriptions: $10 copay - Deductible: $1500 February - $2300 hospital treatment March - Doctor 1: billed $657, she pays $25, after deductible [+$25] - Doctor 2: billed $134, she pays $25, after deductible [+$25] - Doctor 3: billed $232, she pays $25, after deductible [+$25] - Prescription: cost $12, $10 copay [+$10] - Prescription: cost $27, $10 copay [+$10] As we don't consider the February, given we are looking at the share of cost for March, we can see that she visits 3 doctors (\(3\times 25\)), and gets 2 prescriptions (\(2\times 10\)). Therefore we can conclude the amount she pays is as follows: \(3(25)+2(10)\) I will let you solve that simple math problem to find out how much she pays in all. :)
I think this is correct. like he said, it is just some simple math

ykbatman:

um

ykbatman:

thats the correct answer

ykbatman:

but im pretty sure there is an easier way to find it..

grimreaper124:

@ykbatman wrote:
but im pretty sure there is an easier way to find it..
same